Three consecutive odd integers where the sum of the first two is the same as 21 more than the Third
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

integers are x, x+2, and x+4
x+x+2=x+4+21
2x+2=x+25
x=23
x+2=25
x+4=27
Sum of first two is 48 and that is 21 more than 27.

Answer by ikleyn(52799)